OpenStudy (anonymous):

Help? The width of a rectangle is 7 less than its length. If the area of the rectangle is 198, find the perimeter of the rectangle. A 29 B 30 C 58 D 86

OpenStudy (ddcamp):

A = l * w w = l - 7 A = l*(l-7) 198 = l²-7l Solve for l, and then find w. P = 2l + 2w

OpenStudy (anonymous):

Thanks!